Main Street Ice Cream Company uses a plantwide allocation method to allocate overhead based on direct labor-hours at a rate of $3 per labor-hour. Strawberry and vanilla flavors are produced in Department SV. Chocolate is produced in Department C. Sven manages Department SV and Charlene manages Department C. The product costs (per thousand gallons) follow.

 


StrawberryVanillaChocolate
Direct labor (per 1,000 gallons)$750
$825
$1,125
Raw materials (per 1,000 gallons)
800

500

600

 

Required:

a. If the number of hours of labor per 1,000 gallons is 50 for strawberry, 55 for vanilla, and 75 for chocolate, compute the total cost of 1,000 gallons of each flavor using plantwide allocation.

b. Charlene’s department uses older, outdated machines. She believes that her department is being allocated some of the overhead of Department SV, which recently bought state-of-the-art machines. After she requested that overhead costs be broken down by department, the following information was discovered:
 


Department SVDepartment C
Overhead$105,840
$23,760
Machine-hours
25,200

36,000
Labor-hours
25,200

18,000

 
Using machine-hours as the department allocation base for Department SV and labor-hours as the department allocation base for Department C, compute the allocation rate for each.

c. Compute the cost of 1,000 gallons of each flavor of ice cream using the department allocation rates computed in requirement (b) if the number of machine-hours for 1,000 gallons of each of the three flavors of ice cream are as follows: strawberry, 50; vanilla, 55; and chocolate, 150. Direct labor-hours by product remain the same as in requirement (a).
Answer #1

A) step 1) strawberry 50* 3= 150

total cost 750+ 800+ 150= 1,700 answer


vanilla 55* 3= 165

total cost 825+ 500+ 165= 1,490 answer


chocolate 75* 3= 225

total cost 1,125+ 600+ 225= 1,950 answer


B) dept SV 105,840/ 25,200= 4.2 per machine hour answer

Dept C 23,760/ 18,000= 1.32 per labor hour answer


C) strawberry 50* 4.2= 210

total cost 750+ 800+ 210= 1,760 answer


vanilla 55* 4.2= 231

total cost 825+ 500+ 231= 1,556


chocolate step 1) 150-75= 75        step 2) 75* 1.32= 99

1,125+ 600+ 99= 1,824 answer